OpenStudy (alyssajobug):

A baseball is thrown with a vertical velocity of 50 ft/s from an initial height of 6 ft. The height h in feet of the baseball can be modeled by h(t) = -16t^2 + 50t + 6, where t is the time in seconds since the ball was thrown. It takes the ball approximately _____ seconds to reach its maximum height. (Round to the nearest tenth of a second and enter only the number.) To the nearest foot, what is the maximum height that the ball reaches? (Enter only the number.) A player hits a foul ball with an initial vertical velocity of 70 ft/s and an initial height of 5 feet. The maximum height reached by the ball is _____ feet. (Round to the nearest foot and enter only the number.)

OpenStudy (anonymous):

If you are not going to plot this, you can solve it by setting the derivative equal to zero. This is equivalent to finding the time when the velocity v = 50 - 32 t = 0. That time will give you max height when put into h(t).

OpenStudy (loser66):

@douglaswinslowcooper can we solve it by finding the vertex of h(t) ? It is an downward parabola, so the vertex will have t = -b/2a ?

OpenStudy (anonymous):

Sounds right and this agrees with mine.

OpenStudy (loser66):

hihihi, that's good.

OpenStudy (alyssajobug):

I got vertex at 45 when I graphed

OpenStudy (loser66):

nope, it is 1.38888888888888

OpenStudy (loser66):

for t, hehehe

OpenStudy (loser66):

and 45 for h(t) but your question is finding the time, right?

OpenStudy (alyssajobug):

ya

OpenStudy (loser66):

ok, the second part is the height

OpenStudy (alyssajobug):

ok cool ya

OpenStudy (alyssajobug):

height is 45 for second part and isnt 3rd part 82?